# Simulating the economic impact of rationality through reinforcement learning and agent-based modelling

### v1

@@ Abstract
Agent-based models (ABMs) are simulation models used in economics to overcome some of the limitations of traditional frameworks based on general equilibrium assumptions. However, agents within an ABM follow predetermined, not fully rational, behavioural rules which can be cumbersome to design and difficult to justify. Here we leverage multi-agent reinforcement learning (RL) to expand the capabilities of ABMs with the introduction of fully rational agents that learn their policy by interacting with the environment and maximising a reward function. Specifically, we propose a ‘Rational macro ABM’ (R-MABM) framework by extending a paradigmatic macro ABM from the economic literature. We show that gradually substituting ABM firms in the model with RL agents, trained to maximise profits, allows for a thorough study of the impact of rationality on the economy. We find that RL agents spontaneously learn three distinct strategies for maximising profits, with the optimal strategy depending on the level of market competition and rationality. We also find that RL agents with independent policies, and without the ability to communicate with each other, spontaneously learn to segregate into different strategic groups, thus increasing market power and overall profits. Finally, we find that a higher degree of rationality in the economy always improves the macroeconomic environment as measured by total output, depending on the specific rational policy, this can come at the cost of higher instability. Our R-MABM framework is general, it allows for stable multi-agent learning, and represents a principled and robust direction to extend existing economic simulators.
[…]
578
1 Introduction
Traditional economic simulation models are mostly based on the dynamic stochastic general equilibrium (DSGE) framework, and are far from being a perfect representation of real economies, due to their strong assumptions, which typically do not allow for agent heterogeneity, bounded rationality or nonequilibrium dynamics [14, 12].
In today’s tech age, advanced computational methods hold the promise of solving some of these key limitations [25, 8]. Specifically, multi-agent systems (MAS) [33] may offer a bottom-up approach to model economic systems that naturally addresses the above mentioned limitations of traditional models. Since the early 2010s, MAS have gradually been adopted and gained popularity in economics under the name of “agent-based models” (ABMs) [14, 12, 6]. ABMs can model economic systems by simulating a number of individual economic agents, representing decision makers and institutions, which can be defined to be heterogeneous and act with bounded rationality in non-equilibrium environments [29, 28]. Agents typically rely on a pool of hand-crafted behavioural rules, that are defined in advance by economists to mimic the complexity of real economies. While this approach allows addressing some of the limitations described, it also poses new challenges, such as the difficulty in modelling realistic agents, and the complete absence of learning and of ‘rational expectations’ in the agents’ behaviour.
Reinforcement Learning (RL) [27] can offer a solution to classical ABMs limitations as, in principle, it can allow for a grounded introduction of rational expectation and learning in simulation models, without the need of the over-simplistic assumptions of traditional economic modelling frameworks. Moreover, RL can facilitate the design of accurate economic ABMs, as it can eliminate the burden of programming detailed behavioural rules in favour of the easier task of choosing a meaningful reward function, typically known as ‘utility function’ in the economic literature. The recent breakthroughs in deep-RL and multi-agent RL (MARL) [17, 1] along with the ever-growing availability of computational power, are now allowing RL and MARL to emerge as attractive powerful tools to expand the capabilities of traditional ABMs and pave the way to a new generation of economic simulators [10, 2]. Table 1 summarises the main advantages of combining RL with the ABM methodology.
GE ABM ABM+RL
Heterogeneity $\sim$
Rational Expectations $\times$
Bounded Rationality $\times$
Imperfect Information $\times$
Imperfect Markets $\times$
Learning $\times$ $\times$
Ease of design $\times$ $\sim$
Table 1: A simplified overview of the advantages of extending ABMs with RL. While ABMs are more flexible than traditional models based on general equilibrium assumptions, they typically lack a framework for learning rational behaviour, which also increases modelling complexity. Multi-agent RL has the potential of providing this feature in existing ABMs, thus improving their capabilities and facilitating model design at the same time.
Existing work has investigated the use of RL towards economic modelling of different environments [9, 19, 21, 4, 16, 18]. A recent example is the work of Johanson et al. [20], showing that basic phenomena of microeconomics can emerge spontaneously from learning agents in a realistic environment: heterogeneous agents learn to produce, trade and consume resources based on their preferences. Other works focus on real business cycle (RBC) models showing that optimal policies can be learnt through MARL [34, 11, 24, 22, 5, 13], using different learning techniques including curriculum learning [7]. In [24], Mi et al. showcase the advantages of MARL systems over traditional models in the study the impact of taxation policies. Finally, other lines of work focus more on algorithmic pricing  [9] and financial markets simulations [30].
@@ 1 Introduction
Figure 1: The R-MABM model. The left panel shows a schematic diagram of the basis model, in which 4 types of agents (green ovals) exchange goods (yellow rectangles). The arrows represent the flow of the specific good, from provider to receiver. The middle panel illustrates the extension of the basis model we implement with our R-MABM framework. Consumption-good producing firms ‘C-firms’ in the standard model are exclusively ‘bounded rational’ and ‘trend-follower’ and take decisions using simple heuristic rules. These are augmented with ‘fully rational’ RL agents that take decisions in order to maximise profits. The right panel shows a typical learning curve where fully rational RL agents learn to accumulate higher profits than bounded rational agents as the number of learning episodes progresses.
@@
In this paper, we contribute towards this emerging interdisciplinary challenge by proposing the ‘Rational macro ABM’ (R-MABM) framework. The R-MABM framework leverages multi-agent RL to expand the capabilities of classical macro ABMs with the introduction of fully rational agents, which learn their policy by interacting with the environment and maximising their rewards. Specifically, we extend and study the well-known macroeconomic ABM with capital and credit (‘CC-MABM’) from Assenza et al. [3]. Such a model considers an economy populated by households, C-firms, K-firm, and banks. Households supply labour to firms, consume goods, and save money through deposits at the banks. Banks provide loans to firms for investment purposes. K-firms produce capital goods which are then used by C-firms to produce consumption goods. Our extended model substitutes a variable number of C-firms with RL agents, which rationally maximise profits, allowing the study of the impact of rationality in the economy. Figure 1 shows a visualisation of the R-MAMB model, and a sample of the learning process of rational firms.
@@
We experimentally evaluate our framework with extensive simulations, considering different levels of market competition (controlled by the number of firms visible to each consumer), rationality (controlled by the number of RL agents), and agent learning with shared or independent policies. We show that RL agents spontaneously learn three distinct strategies for maximising profits, with the optimal strategy depending on the level of market competition and rationality. We also find that agents with independent policies spontaneously learn to segregate into different strategic groups, thus increasing market power and overall profits, without explicitly communication with each other. Finally, we assess the macroeconomic impact of higher degrees of rationality by measuring the direct effects on GDP and on its volatility. To the best of our knowledge, our R-MABM framework is the first attempt at extending macroeconomic ABMs with the introduction of rational agents using RL. Our results amply demonstrate the robustness of the R-MABM framework and its relevance to a vast community at the intersection between economics and computer science, thus paving the way to numerous applications and future investigations.
2 The Rational-MABM model
@@
2.1 The basis model
We here briefly describe the basis macro ABM, namely the macroeconomic ABM with capital and credit (CC-MABM) from Assenza et al [3]. For a more detailed description of the model and its parameters we refer to the original manuscript.
Agent types.
The basis model comprises three types of agents defined as follows.
  • Households consist of workers and capitalists. Workers supply labour, receiving a wage. Capitalists receive dividends and do not work. Both workers and capitalists deposit accumulated wealth in banks.
  • Banks receive deposits from households and extend loans to firms.
@@ 2.1 The basis model
Markets.
All of the households aim to buy consumption goods, and therefore participate in a search and matching consumption market, defined as follows.
@@
A similar mechanism is in place within the capital goods market: each C-firm randomly visits $z_{k}$ K-firms and buys the required amount of capital goods starting from the firm offering the lowest price. Note that, a higher (lower) value of the $z_{k}$ and $z_{c}$ parameters, implies that each customer will be able to compare the prices of a higher (lower) number of firms. Hence, importantly, we can consider those two parameters as effective controllers of the level of competition of the search and matching markets described.
Price and quantity decisions of firms.
The model assumes that at each time step $t$, a firm $i$ decides a price $P_{i,t}$ at which to sell its goods, and a target quantity $Y*_{i,t}$ of goods it aims to produce. Based on the desired production $Y*_{i,t}$, the firm decides how many workers to employ $N_{i,t}$ and how much capital $K_{i,t}$ is required. The firm will then attempt to acquire labour and capital in the corresponding markets. Finally, the production function follows a Leontief technology, i.e., $Y_{i,t}=\min\ (\alpha_{N}N_{i,t},\alpha_{K}K_{i,t})$ where $\alpha_{N}$ and $\alpha_{K}$ are labour and capital productivity, respectively and $Y_{i,t}$ is the quantity of goods actually produced. Notice that consumption goods are non-storable, so unsold goods do not carry over to the next period.
For the rest of this work, we will often call a defined policy to set $P_{i,t}$ and $Y*_{i,t}$ a ‘price-quantity strategy’, or simply a ‘strategy’. The price-quantity strategy of C-firms in the basis model is based on the following market signals. At the end of the period $t$, the firm $i$ observes the amount of goods sold $Y^{s}_{i,t}=\min\left(Y_{i,t},Y_{i,t}^{d}\right)$, where $Y_{i,t}^{d}$ is the actual demand, and the average price $P_{t}$ charged by competitors.
@@
$P_{i,t+1}$ =$=\left\{\!\!\begin{array}[]{ll}P_{i,t}(1+\eta_{i,t+1})&\text{ if }\Delta_{i,t}^{Y}\leq 0\land\Delta_{i,t}^{P}<0\\[5.69054pt] P_{i,t}(1-\eta_{i,t+1})&\text{ if }\Delta_{i,t}^{Y}>0\land\Delta_{i,t}^{P}\geq 0\\[5.69054pt] P_{i,t}&\!\text{ otherwise}\end{array}\right.$ (4)
$Y^{*}_{i,t+1}$ =$=\left\{\!\!\begin{array}[]{ll}Y_{i,t}+\rho|\Delta_{i,t}^{Y}|&\text{ if }\Delta_{i,t}^{Y}\leq 0\land\Delta_{i,t}^{P}\geq 0\\[5.69054pt] Y_{i,t}-\rho|\Delta_{i,t}^{Y}|&\text{ if }\Delta_{i,t}^{Y}>0\land\Delta_{i,t}^{P}<0\\[5.69054pt] Y^{*}_{i,t}&\!\text{ otherwise}\end{array}\right.,$ (8)
@@ 2.2 The RL agents
We expand the basis model by introducing fully rational C-firms as RL agents that learn their price-quantity strategies by interacting with the environment and maximising a reward function. In particular, the underlying decision model for a subset $N$ of fully rational C-firms corresponds to a stochastic game (see e.g. [26, 31]) with $N$ players, i.e. the RL agents, or, in the special case when $N=1$, a Markov decision process. The remaining bounded rational agents, those of the basis model, are not seen as players of the game since they all act following the same fixed behavioural strategy, and so their actions can just be seen as part of the probabilistic transition of the stochastic game.
The aim of each RL agent is to maximise the sum of its discounted rewards
$S_{i}=\sum_{t=1}^{t_{\textit{sim}}}\gamma^{t}r_{i,t}.$ (9)
@@
The maximisation of rewards (to be later defined as firms profits) is the feature of the RL agents that allows them to be fully rational, in contrast with standard firms of the basis model that simply follow a behavioural trend-following heuristic.
@@
We train each RL agent using a standard asynchronous Q-learning algorithm, like that introduced in [32], with $\epsilon$ greedy policy. Therefore, each RL agent considers discrete state and action spaces, namely $\mathcal{S}$ and $\mathcal{A}$, and learns a value function $Q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, through the well-known Bellman equation, to estimate the expected reward $r$ for each state-action pair $(s,a)$. Note that, as explained below, the agents do not have full knowledge of the state of the model, nor can they fully observe the actions performed by the other agents. Because of this, we resort to such a technique, where each agent regards the others as part of the environment.
State space.
@@
Similarly to standard (bounded rational) firms, RL agents use the price delta $\tilde{\Delta}^{P}_{i,t}$ and the firm stock $\tilde{\Delta}^{Y}_{i,t}$ values as state observations $\mathcal{S}$. This means that the current state of the underlying stochastic game model is only partially observable by agents. In the case of RL agents, the two quantities are defined by first taking the logarithms of the corresponding quantities
$\tilde{\Delta}^{P}_{i,t}$ $=\log(P_{i,t})-\log(P_{t})=\log(P_{i,t}/P_{t})$ (10)
$\tilde{\Delta}^{Y}_{i,t}$ $=\log(Y_{i,t})-\log(Y_{i,t}^{d})=\log(Y_{i,t}/Y_{i,t}^{d}).$ (11)
@@
Since tabular Q-learning requires a finite observation space, these quantities are binned. To this end, a minimum and maximum value are set, $\mathcal{S}_{min}$ and $\mathcal{S}_{max}$ respectively. Then, the interval delimited by such values is uniformly divided in $n_{\mathcal{S}}$ poles. Each real-valued observation is then assigned the index of the nearest pole. In this way, the observation space is discretised over a finite number of equally spaced bins. The logarithmic transformation in the state space definition above allows us to interpret the two states as the percentage differences of the past choices with respect to past market signals.
Action space.
@@
Based on the current state observations, each RL agent $i$ can make price-quantity decisions for the next step, thus setting $P_{i,t+1}$ and $Y^{*}_{i,t+1}$, respectively. To change these values the agent selects an action set $a_{i,t}=(a_{i,t}^{Y},a_{i,t}^{P})$. Given the selected actions, prices and quantities are adjusted as
$P_{i,t+1}$ $=e^{\log(P_{i,t})+a_{i,t}^{P}}$ (12)
$Y^{*}_{i,t+1}$ $=e^{\log(Y_{i,t}^{*})+a_{i,t}^{Y}}.$ (13)
@@
Reward.
At each step $t$, each RL agent observes $s_{i,t}=\left(\tilde{\Delta}^{P}_{i,t},\tilde{\Delta}^{Y}_{i,t},\right)$ performs action $a_{i,t}=\left(a_{i,t}^{P},a_{i,t}^{Y}\right)$, and obtains a reward $r_{i,t}$ given by
$r_{i,t}=\left\{\!\!\begin{array}[]{ll}\pi_{i,t}&\text{ if }A_{i,t}>0\\[5.69054pt] -100&\text{ otherwise}\end{array}\right.,$ (14)
@@
From a game-theoretical standpoint, one could also think of determining the underlying stochastic game and compute the corresponding optimal policies for the RL agents, given some solution concept such as Markov perfect equilibrium [23]. While this could theoretically be done as the basis model can provide a full definition of the states, actions, transition and rewards, it would not be feasible in practice as the size of the model and of the strategy space are highly prohibitive. Moreover, the possible use of information extracted from the game model would not represent a realistic assumption since, in the real world, agents are not able to fully observe their environment. On the other hand, the adopted Q-learning procedure allows each agent to learn a policy which approximates the best response for the behaviour of bounded rational agents and, possibly, the other RL agents as well.
Training.
During the training each agent $i$ updates the Q-function according to the following Bellman equation
@@
During the training phase each agent follows an $\epsilon$-greedy action selection
$a_{i,t}=\left\{\!\!\begin{array}[]{ll}\text{random action}&\text{ with probability }\epsilon\\[2.84526pt] \arg\max\limits_{a}Q(s_{i,t},a)&\text{ otherwise}\end{array}\right.,$ (16)
@@
$\rho$ Quantity adjustment parameter $0.9$
$\bar{\eta}$ Price adjustment parameter 0.1
@@
$\mathcal{S}_{min}$ Observation lower bound $-2$
$\mathcal{S}_{max}$ Observation upper bound $2$
$n_{\mathcal{S}}$ Number of observations $11$
$\mathcal{A}_{min}$ Actions lower bound $-0.3$
$\mathcal{A}_{max}$ Actions upper bound $0.3$
$n_{\mathcal{A}}$ Number of discrete actions $11$
@@
Table 2: Experimental parameters used. List of parameters used in the simulations. The values of $z_{c}$ and $N$ range from $2$ to $10$ and from $1$ to $20$ respectively, the specific values used can be read directly in the figures of Section 4.
3 Experimental setup
We simulate the macroeconomic model considering a single bank, $1000$ workers, $20$ K-firms, and $100$ C-firms. In our experiments, we vary the levels of competition (i.e., $z_{c}$) from $2$ to $10$, and the number $N$ of RL agents, representing fully rational C-firms, from $1$ to $20$. Since we focus exclusively on C-firms in our experiment, we will refer to them simply as ‘firms’ for the remainder of the work. All simulation parameters are described in Table 2, while additional parameters of the basis model are those of Assenza et al. [3].
We train the RL agents for each model configuration, considering both shared and independent policies. We consider $T_{\textit{train}}=100$ episodes for the training phase, while the test phase consists of $T_{\textit{test}}=20$ episodes. Unless otherwise specified, we report averages of the simulation results over the 20 testing episodes, with error bars representing one standard deviation.
Similarly to some existing MARL frameworks for economic modelling, we consider a sort of curriculum learning [7]: the simulation starts without the RL agents, which are introduced after $t_{\textit{burn-in}}$ steps. Then, the RL agents assume control of $N$ firms for additional $t_{\textit{sim}}$ steps. At the start of the training phase, RL agents are initialised with an empty Q-matrix i.e. all values are zero. During training, and at each episode $\tau=0,\dots,T_{\textit{train}}-1$ , the agents update the exploration rate $\epsilon$ of the $\epsilon$-greedy policy in Eq. (16) as $\epsilon\leftarrow\max\left(0.9^{\tau},0.01\right)$. Simulation results are presented in the next section.
4 Results
In this section we discuss the macroeconomic simulations obtained with our R-MABM framework, where we gradually substitute standard firms with RL agents trained to maximise profits. In particular, we showcase how our framework allows for a detailed study of the impact of rationality in a complex economic system. We divide our results into microeconomic effects (Section 4.1) and macroeconomic effects (Section 4.2). To the best of our knowledge, our framework is the first to allow for similar analyses.

### v2

@@ Abstract.
Agent-based models (ABMs) are simulation models used in economics to overcome some of the limitations of traditional frameworks based on general equilibrium assumptions. However, agents within an ABM follow predetermined ‘bounded rational’ behavioural rules which can be cumbersome to design and difficult to justify. Here we leverage multi-agent reinforcement learning (RL) to expand the capabilities of ABMs with the introduction of ‘fully rational’ agents that learn their policy by interacting with the environment and maximising a reward function. Specifically, we propose a ‘Rational macro ABM’ (R-MABM) framework by extending a paradigmatic macro ABM from the economic literature. We show that gradually substituting ABM firms in the model with RL agents, trained to maximise profits, allows for studying the impact of rationality on the economy. We find that RL agents spontaneously learn three distinct strategies for maximising profits, with the optimal strategy depending on the level of market competition and rationality. We also find that RL agents with independent policies, and without the ability to communicate with each other, spontaneously learn to segregate into different strategic groups, thus increasing market power and overall profits. Finally, we find that a higher number of rational (RL) agents in the economy always improves the macroeconomic environment as measured by total output. Depending on the specific rational policy, this can come at the cost of higher instability. Our R-MABM framework allows for stable multi-agent learning, is available in open source, and represents a principled and robust direction to extend economic simulators.
agent-based modelling, reinforcement learning, macroeconomics
ccs: Applied computing Economicsccs: Computing methodologies Modeling and simulation Simulation types and techniques Agent / discrete modelsccs: Artificial intelligence Planning and scheduling Planning under uncertaintyccs: Machine learning Machine learning approaches Markov decision processes
@@
Figure 1. The R-MABM model. The left panel shows a schematic diagram of the basis model, in which 4 types of agents (green ovals) exchange goods (yellow rectangles). The arrows represent the flow of the specific good, from provider to receiver. The middle panel illustrates the extension of the basis model we implement with our R-MABM framework. Consumption-good producing firms ‘C-firms’ in the standard model are exclusively ‘bounded rational’ and take decisions using a heuristic trend-following rule. These are augmented with ‘fully rational’ RL agents that take decisions in order to maximise profits. The right panel shows a typical learning curve where fully rational RL agents learn to accumulate higher profits than bounded rational agents as the number of learning episodes progresses.
1. Introduction
Traditional economic simulation models are mostly based on general equilibrium frameworks, and are far from being a perfect representation of real economies, due to their strong assumptions, which typically do not allow for agent heterogeneity, bounded rationality or nonequilibrium dynamics (Farmer and Foley, 2009; Dawid and Gatti, 2018). Today, advanced computational methods hold the promise of solving some of these key limitations (Dawid et al., 2008; Rothe et al., 2015; Blume et al., 2015; Benedetti et al., 2022; Monti et al., 2023). Specifically, multi-agent systems (MAS) (Wooldridge, 2009) may offer a bottom-up approach to model economic systems that naturally addresses the above mentioned limitations of traditional models. Since the early 2010s, MAS have gradually been adopted and gained popularity in economics under the name of “agent-based models” (ABMs) (Farmer and Foley, 2009; Dawid and Gatti, 2018; Axtell and Farmer, 2022). ABMs can model economic systems by simulating a number of individual economic agents, representing decision makers and institutions, which can be defined to be heterogeneous and act with bounded rationality in non-equilibrium environments (Tesfatsion, 2023, 2006). These agents typically rely on a pool of hand-crafted behavioural rules, that are defined in advance by economists to mimic the complexity of real economies (Vyetrenko et al., 2020). While this approach allows addressing some of the limitations described, it also poses new challenges, such as the difficulty in defining realistic behaviours, and the absence of learning agents and of ‘rational’ agents that optimise their behaviour based on a specific objective.
Reinforcement Learning (RL) (Sutton and Barto, 2018) can offer a solution to classical ABM limitations as, in principle, it can allow for a grounded introduction of rational optimisation and learning in simulation models, without the need for the over-simplistic assumptions of traditional economic modelling frameworks. Moreover, RL can facilitate the design of accurate economic ABMs, as it can eliminate the burden of programming detailed behavioural rules in favour of the easier task of choosing meaningful reward functions, typically known as ‘utility functions’ in the economic literature.
@@ 1. Introduction
Our contribution. In this work, we contribute towards this interdisciplinary challenge by proposing a framework to leverage multi-agent RL to expand the capabilities of classical macro ABMs. Specifically, we extend and study the well-known macroeconomic ABM with capital and credit (‘CC-MABM’) from Assenza et al. (Assenza et al., 2015). Such a model considers an economy populated by households, firms and banks. Households supply labour to firms, consume goods, and save money through deposits at the banks, and banks provide loans to firms. Our extended model substitutes a variable number of firms with RL agents. Importantly, we refer to the firms of the original models as ‘bounded rational’ in the sense that they behave according to fixed rules that are not the result of a maximisation of a utility function, while we refer to RL firms as ‘fully rational’ since they act in order to explicitly maximise profits. From these considerations, we name our framework ‘Rational macro ABM’ (R-MAMB). Figure 1 shows a visualisation of the R-MAMB model, and a sample of the learning process of rational firms.
@@
We experimentally evaluate our framework with extensive simulations, and show that RL agents spontaneously learn three distinct strategies for maximising profits, with the optimal strategy depending on the level of market competition and rationality in the economy. We also find that agents with independent policies spontaneously learn to segregate into different strategic groups, thus increasing market power and overall profits, without explicitly communicating with each other. Finally, we assess the macroeconomic impact of higher degrees of rationality by measuring the direct effects on GDP and on its volatility. To the best of our knowledge, R-MABM is the first attempt at extending macroeconomic ABMs with the introduction of rational agents using RL. Our results amply demonstrate the robustness of the R-MABM framework and its relevance to a vast community at the intersection between economics and computer science, thus paving the way for numerous applications and future investigations. Accordingly, we made the code for the R-MABM easy to use and freely available  [^1] The R-MABM is available in open source at https://github.com/Brusa99/R-MABM, while the code used for the basis model is available at https://github.com/bancaditalia/ABCredit.jl..
Related work. Thanks to recent breakthroughs in deep-RL and multi-agent RL (MARL) (Gronauer and Diepold, 2022; Albrecht et al., 2023), along with the ever-growing availability of computational power, RL and MARL are rapidly emerging as attractive powerful tools to expand the capabilities of traditional ABMs and pave the way to a new generation of economic simulators (Charpentier et al., 2021; Ardon et al., 2023; Dwarakanath et al., 2024). Several works have investigated the use of RL towards economic modelling for different environments and settings (Atashbar and Shi, 2022), including learning monetary policies  (Hinterlang and Tänzer, 2021), calibrating economic ABMs (Glielmo et al., 2023), solving heterogeneous general equilibrium models (Hill et al., 2021), approaching consumption-saving problems (Kuriksha, 2021), studying social segregation dynamics (Sert et al., 2020) or mortgage relief strategies (Garg et al., 2024), modelling agents with heterogeneous bounded rationality (Evans and Ganesh, 2024), or modelling decision making process in technology uptake models (Klügl and Kyvik Nordås, 2023). Another very active line of research focuses on the use of RL in the simulation of financial stock markets  (Vadori et al., 2024; Coletta et al., 2022; Albers et al., 2024).
For models of the entire economy, a recent example is the work of Johanson et al. (Johanson et al., 2022), which shows that basic economic phenomena of microeconomics can emerge spontaneously from learning agents. Other work focuses on real business cycle (RBC) models showing that optimal policies can be learnt through MARL (Zheng et al., 2022; Curry et al., 2023; Mi et al., 2023; Lindau, 2023; Atashbar and Shi, 2023), using different learning techniques including curriculum learning (Bengio et al., 2009). In (Mi et al., 2023), Mi et al. showcase the advantages of MARL systems over traditional models in designing and studying taxation policies.
As we focus on RL modelling of firms in our macro ABM for this study, our work is also related to the vast literature on algorithmic pricing using RL, pioneered by Tesauro and Kephart (Tesauro and Kephart, 1998, 2002) and more recently taken up by many studies in both economics and computer science (Waltman and Kaymak, 2008; Cai et al., 2018; Calvano et al., 2020; Kastius and Schlosser, 2022).
Finally, our work is related to (Dosi et al., 2020), where Dosi et al. study the effects of substituting the traditional bounded rational behavioural rules of a macroeconomic ABM with more complex adaptive rules, without using RL.
2. The Rational-MABM model
@@
2.1. The basis model
We here briefly describe the basis macro ABM, namely the macroeconomic ABM with capital and credit (CC-MABM) from Assenza et al (Assenza et al., 2015). For a more detailed description of the model and its parameters we refer to the original manuscript.
Agent types. The basis model comprises three types of agents defined as follows.
  • Households consist of workers and capitalists. Workers supply labour, receiving a wage. Capitalists receive dividends and do not work. Both deposit accumulated wealth in banks.
  • Banks receive household deposits and extend loans to firms.
@@ 2.1. The basis model
Markets. All of the households aim to buy consumption goods, and therefore participate in a search and matching consumption market, defined as follows.
@@
Note that, a higher (lower) value of $z_{c}$ implies that each consumer will be able to compare the prices of a higher (lower) number of firms. Hence, importantly, we can consider those this parameter as an effective controller of the level of competition of the search and matching market described.
Price and quantity decisions of firms. The model assumes that at each time step $t$, a firm $i$ decides a price $P_{i,t}$ at which to sell its goods, and a target quantity $Y^{*}_{i,t}$ of goods it aims to produce. Based on the desired production $Y^{*}_{i,t}$, the firm decides how many workers to employ $N_{i,t}$ and how much capital $K_{i,t}$ is required. The firm will then attempt to acquire labour and capital in the corresponding markets. Finally, the production function follows a Leontief technology, i.e., $Y_{i,t}=\min\ (\alpha_{N}N_{i,t},\alpha_{K}K_{i,t})$ where $\alpha_{N}$ and $\alpha_{K}$ are labour and capital productivity, respectively and $Y_{i,t}$ is the quantity of goods actually produced. Notice that consumption goods are non-storable, so unsold goods do not carry over to the next period.
For the rest of this work, we will often call a defined policy to set $P_{i,t}$ and $Y^{*}_{i,t}$ a ‘price-quantity strategy’, or simply a ‘strategy’. The price-quantity strategy of C-firms in the basis model is based on the following. At the end of the period $t$, the firm $i$ observes the amount of goods sold $Y^{s}_{i,t}=\min\left(Y_{i,t},Y_{i,t}^{d}\right)$, where $Y_{i,t}^{d}$ is the actual demand, and the average price $P_{t}$ charged by competitors.
@@
(4) $P_{i,t+1}$ =$=\left\{\!\!\begin{array}[]{ll}P_{i,t}(1+\eta_{i,t+1})&\text{ if }\Delta_{i,t}^{Y}\leq 0\land\Delta_{i,t}^{P}<0\\[5.69054pt] P_{i,t}(1-\eta_{i,t+1})&\text{ if }\Delta_{i,t}^{Y}>0\land\Delta_{i,t}^{P}\geq 0\\[5.69054pt] P_{i,t}&\!\text{ otherwise}\end{array}\right.$
(8) $Y^{*}_{i,t+1}$ =$=\left\{\!\!\begin{array}[]{ll}Y_{i,t}+\rho|\Delta_{i,t}^{Y}|&\text{ if }\Delta_{i,t}^{Y}\leq 0\land\Delta_{i,t}^{P}\geq 0\\[5.69054pt] Y_{i,t}-\rho|\Delta_{i,t}^{Y}|&\text{ if }\Delta_{i,t}^{Y}>0\land\Delta_{i,t}^{P}<0\\[5.69054pt] Y^{*}_{i,t}&\!\text{ otherwise}\end{array}\right.,$
@@ 2.2. The RL agents
We expand the basis model by introducing fully rational C-firms as RL agents that learn their price-quantity strategies by interacting with the environment and maximising a reward function. In particular, the underlying decision model for a subset $N$ of fully rational C-firms corresponds to a stochastic game (see e.g. (Shapley, 1953; van der Wal, 1981)) with $N$ players, i.e. the RL agents, or, in the special case when $N=1$, a Markov decision process. The remaining bounded rational agents, those of the basis model, are not players of the game as they all act following the same fixed behavioural rules, and they can hence be seen as part of the probabilistic transition of the stochastic game.
Each RL agent aims to maximise the sum of its discounted rewards
(9) $S_{i}=\sum_{t=1}^{t_{\textit{sim}}}\gamma^{t}r_{i,t}.$
@@
The maximisation of rewards (to be later defined as firms’ profits) is the feature of the RL agents by which we refer to them as fully rational, in contrast with bounded rational firms that follow a behavioural heuristic that is not the result of an explicit optimisation.
@@
We train each RL agent using a standard asynchronous Q-learning algorithm, like that introduced in (Watkins and Dayan, 1992), with $\epsilon$ greedy policy. Therefore, each RL agent considers discrete state and action spaces, namely $\mathcal{S}$ and $\mathcal{A}$, and learns a value function $Q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, through the well-known Bellman equation, to estimate the expected reward $r$ for each state-action pair $(s,a)$. Note that, as explained below, the agents do not have full knowledge of the state of the model, nor can they fully observe the actions performed by the other agents. Because of this, we resort to such a technique, where each agent regards the others as part of the environment.
@@
State space. Similarly to standard (bounded rational) firms, RL agents use the price delta $\tilde{\Delta}^{P}_{i,t}$ and the firm stock $\tilde{\Delta}^{Y}_{i,t}$ values as state observations $\mathcal{S}$. This means that the current state of the underlying stochastic game model is only partially observable by agents. In the case of RL agents, the two quantities are defined by first taking the logarithms of the corresponding quantities
(10) $\tilde{\Delta}^{P}_{i,t}$ $=\log(P_{i,t})-\log(P_{t})=\log(P_{i,t}/P_{t})$
(11) $\tilde{\Delta}^{Y}_{i,t}$ $=\log(Y_{i,t})-\log(Y_{i,t}^{d})=\log(Y_{i,t}/Y_{i,t}^{d}).$
@@
Since tabular Q-learning requires a finite observation space, these quantities are binned in $n_{\mathcal{S}}$ equally spaced poles from a minimum $\mathcal{S}_{min}$ to a maximum $\mathcal{S}_{max}$. Each real-valued observation is then assigned the index of the nearest pole. The logarithmic transformation in the state space definition above allows us to interpret the two states as the percentage differences of the past choices with respect to past market signals.
@@
Action space. Based on the current state observations, each RL agent $i$ can make price-quantity decisions for the next step, thus setting $P_{i,t+1}$ and $Y^{*}_{i,t+1}$, respectively. To change these values the agent selects an action set $a_{i,t}=(a_{i,t}^{Y},a_{i,t}^{P})$. Given the selected actions, prices and quantities are adjusted as
(12) $P_{i,t+1}=e^{\log(P_{i,t})+a_{i,t}^{P}},\quad\quad\qquad Y^{*}_{i,t+1}=e^{\log(Y_{i,t}^{*})+a_{i,t}^{Y}}.$
@@
Reward. At each step $t$, each RL agent observes $s_{i,t}=\left(\tilde{\Delta}^{P}_{i,t},\tilde{\Delta}^{Y}_{i,t},\right)$ performs action $a_{i,t}=\left(a_{i,t}^{P},a_{i,t}^{Y}\right)$, and obtains a reward $r_{i,t}$ given by
(13) $r_{i,t}=\left\{\!\!\begin{array}[]{ll}\pi_{i,t}&\text{ if }A_{i,t}>0\\[5.69054pt] -100&\text{ otherwise}\end{array}\right.,$
@@
From a game-theoretical standpoint, the underlying stochastic game is guaranteed to admit optimal policies for the RL agents, given some solution concept such as Markov perfect equilibrium (Maskin and Tirole, 2001). However, the corresponding problem is known to be at least PPAD-hard, since finding a Nash equilibrium is already PPAD-complete (see e.g. (Daskalakis et al., 2009)). Therefore, while this could theoretically be done as the basis model can provide a full definition of the states, actions, transition and rewards, it would not be feasible in practice as the size of the model and of the strategy space are highly prohibitive. Moreover, the possible use of information extracted from the game model would not represent a realistic assumption since, in the real world, agents are not able to fully observe their environment. On the other hand, the adopted Q-learning procedure allows each agent to learn a policy which approximates the best response for the behaviour of bounded rational agents and, possibly, the other RL agents as well.
Training. During the training each agent $i$ updates the Q-function according to the following Bellman equation
@@
During training, each agent follows an $\epsilon$-greedy action selection
(15) $a_{i,t}=\left\{\!\!\begin{array}[]{ll}\text{random action}&\text{ with probability }\epsilon\\[2.84526pt] \arg\max\limits_{a}Q(s_{i,t},a)&\text{ otherwise}\end{array}\right.,$
@@
Table 1. Experimental parameters used. List of parameters used in the simulations. The specific values used for $z_{c}$ and $N$ can be read directly in the figures of Section 4.